# Reweight-annealing method for calculating the value of partition function via quantum Monte Carlo

### v3

@@ Appendix A Supplemental material
One important application of our method is to calculate the entanglement entropy in conjunction with the Bisognano-Wichmann theorem, which offers an excellently approximated functional form to the entanglement Hamiltonian in systems exhibiting translational symmetry. We briefly introduce the functional form of 1D/2D system with periodic boundary condition here.
@@
In addition to the LBW-HM model shown in the main body of the paper, we present another example of the 1D transverse field Ising model (TFIM), and calculate its Von Neumann entropy at the critical point with the reweight-annealing method. The exact result is $S(L)\propto(c/3)\ln L$, where $c=0.5$ is the central charge, and our method gives $c\approx 0.501(4)$, as it shown in Fig. 3.

### v5

@@ Appendix B Computing entanglement entropy
One important application of our method is to calculate the entanglement entropy in conjunction with the Bisognano-Wichmann theorem, which offers an excellently approximated functional form to the entanglement Hamiltonian in systems exhibiting translational symmetry [72, 73]. We briefly introduce the functional form of 1D/2D system with periodic boundary condition here.
@@
In addition to the LBW-HM model shown in the main body of the paper, we present an additional example of the 1D transverse field Ising model (TFIM), and calculate its Von Neumann entropy at the critical point with the reweight-annealing method. The exact result is $S(L)\propto(c/3)\ln L$, where $c=0.5$ is the central charge, and our method gives $c\approx 0.501(4)$, as it shown in Fig. 7.